James scored 64% on the first test, and 70% on the second. Was James right in thinking that he had a much better result in the second test?
Term 1 test mean=58% Standard deviation=6
Term 2 mean=65% Standard deviation= 5

Answer by stanbon(75887)

James scored 64% on the first test, and 70% on the second. Was James right in thinking that he had a much better result in the second test?
Term 1 test mean=58% Standard deviation=6
Term 2 mean=65% Standard deviation= 5
-------------
Procedure:: Compare the z-scores
---
z(64) = (64-58)/6 = 7/6
z(70) = (70-65)/5 = 1
-----
Ans: Since 7/6 > 1, the 1st test score was better.
----------
Cheers,
Stan H.